### v1

# An insight on some properties of high order nonstandard linear multistep methods

Abstract
In this paper, nonstandard multistep methods are observed. It is shown that under some (sufficient and necessary) conditions, these methods attain the same order as their standard counterparts - to prove this statement, a nonstandard version of Taylor’s series is constructed. The preservation of some qualitative properties (boundedness, a weak form of monotonicity, and the linear combination of the components) is also proven for all step sizes. The methods are applied to a one-dimensional equation and a system of equations, in which the numerical experiments confirm the theoretical results.
keywords:
nonstandard finite difference, multistep method, positivity preservation, SSP linear multistep methods
pacs:
MSC Classification]65L06, 92D25, 92D30
1 Introduction
Two key aspects of the numerical modeling of differential equations have always been the high accuracy of the methods along with the preservation of important features of the initial continuous problem. It has been known that the application of high-order methods does not necessarily imply the latter (see e.g. [1], along with the specific cases of multistep methods: [2, 3]).
During the previous decades, numerous different approaches have been taken to construct methods with such properties. One such approach, which has had an increasing prevalence in recent years, has been the approach of nonstandard methods: only last year, multiple articles applied these ideas to epidemic problems [4, 5, 6, 7, 8], to equations of population dynamics [9, 10, 11, 12], to models in microbiology and chemistry [13, 14], to problems in finance [15, 16] and also to equations with derivatives of fractional order [17, 18, 19, 20], just to name a few (for a comprehensive list of earlier applications, the author recommends the remarkably insightful survey articles of Patidar [21, 22] and the references therein). The main advantage of the use of such methods is that they are customarily constructed in a way that some desired properties of the original, continuous model are preserved for any choice of step size.
One thing that most former applications of the nonstandard ideas lack is a discussion regarding the order of the methods: the majority of them achieved only the order of one. Nonetheless, in recent years, there has been an emerging trend for the construction of higher-order nonstandard methods.
Earliest examples include the application of these ideas to the Yee algorithm [23, 24] and the observation of the second-order nonstandard Runge-Kutta method [25, 26]. Nevertheless, nowadays one leading school of thought is the use of theta methods, which can achieve an order of two [27, 28], can be combined with second order Runge-Kutta methods [29], and can also be applied to reaction-diffusion [30, 31] and singularly perturbed problems [32].
Another possibility is the application of Richardson extrapolation [33], which can also be applied to population models [34]. There are also approaches involving the use of a more complicated form of the denominator function [35, 36, 37, 38], while other ideas incorporate innovative modifications of the right-hand side of the method [39, 40]. Other works include the general order conditions for one-step nonstandard methods applied to some special types of equations [41, 42] or methods with variable step sizes [43]. These ideas have also been applied to several real-life applications [13, 44, 45, 10].
It should also be mentioned that the application of strongly stable methods in the context of nonstandard methods (which will be the core idea in the preservation of the properties in our case) has appeared in some previous works: some have combined the nonstandard methods with total variation diminishing (TVD) schemes [46] and Runge-Kutta methods written in a strongly stability preserving (SSP) form [47] - the authors in [48] even got the order condition for second order methods akin to the one proven in Section 3.1 (other similar instances include [35] and [38]).
One article to highlight is the work of Dang and Hoang [49] involving nonstandard Runge-Kutta methods. In this, it is shown that the nonstandard versions of these schemes have the same order as their standard counterparts, and (under some conditions) they also preserve the properties of the initial differential equation. It should also be pointed out that the condition of convergence in that work corresponds strongly to the one that is proven in this present article in Section 3.1 (however, their condition is only sufficient while the one in this paper is also necessary).
In this paper, we focus on nonstandard multistep methods. Such schemes were introduced in [50] by Anguelov and Lubuma. However, due to the issues mentioned in Section 2, these ideas were abandoned and not considered by other authors. This work aims to establish that such issues can be resolved by imposing some additional conditions on the denominator function. In Section 2, we state the most important definitions and core concepts. The main results are presented in Section 3 with their proofs detailed in Section 4. Two possible applications are mentioned in Section 5 with several numerical solutions illustrating the effectiveness of the methods. The article concludes with some possible extensions in Section 6.
2 Preliminaries
In this paper, we observe an initial problem of an autonomous ordinary differential equation given in the form
$\begin{cases}y^{\prime}(t)&=f(y(t)),\\ y(t_{0})&=\tilde{y},\end{cases}$ (1)
where $y:[t_{0},\;T]\rightarrow\mathbb{R}^{m}$ ($[t_{0},\;T]\subset\mathbb{R}$, $m\in\mathbb{Z}^{+}$), $y(t)=\left(y_{1}(t),\;y_{2}(t),\;\dots,\;y_{m}(t)\right)^{T}$, $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ and $\tilde{y}\in\mathbb{R}^{m}$. We assume that (1) has a unique solution (this can be ensured if $f$ has the Lipschitz property).
[…]
$D_{\Delta t}u^{n}=F_{\Delta t}(f;u^{n}),$ (2)
where $D_{\Delta t}u^{n}$ is the approximation of the derivative $y^{\prime}(t)$ of the exact solution $y(t)$, and $F_{\Delta t}(f;u^{n})$ approximates the right-hand side of the differential equation. Later, the notation $F_{\Delta t}(u^{n})$ is also used.
The main goal of this paper is to extend the well-known class of explicit linear multi-step methods with $k$-many steps (see e.g. [51]) given in the form
$\sum_{j=0}^{k}\alpha_{j}u^{n+j}=\Delta t\sum_{j=0}^{k-1}\beta_{j}f(u^{n+j}).$ (3)
Subsequently, we will also consider the following form used in the theory of strongly stability preserving methods (see e.g. [52])
$u^{n+1}=\sum_{j=1}^{s}\left(\tilde{\alpha}_{j}u^{n+1-j}+\Delta t\tilde{\beta}_{j}f(u^{n+1-j})\right),$ (4)
which is beneficial when observing the preservation of certain qualitative properties of the method. (We can also include implicit methods if the summation starts at $j=0$.)
In [50], Anguelov and Lubuma proposed the nonstandard version of the previous (’standard’) method (in line with the nonstandard modeling rules of Mickens, see e.g. [1]), which takes the form
$\sum_{j=0}^{k}\alpha_{j}u^{n+j}=\varphi(\Delta t)\sum_{j=0}^{k}\beta_{j}f(u^{n+j}),$ (5)
where $\varphi$ is a nonnegative real-valued one-variable function such that
$\varphi(x)=x+O(x^{2})\qquad\text{as }x\rightarrow 0.$ (6)
In that work, the consistency and the zero-stability of the non-standard method were proven. Namely,
Theorem 1 ([50], Theorem 6).
If (3) represents a standard finite difference scheme that is consistent and zero-stable, then any corresponding nonstandard finite difference scheme (5) (that is, with the same $\alpha_{j}$ and $\beta_{j}$ constants) is necessarily consistent. Furthermore, scheme (5) is zero-stable provided that the operator $F_{\Delta t}$ satisfies, for some $M>0$ independent of $h$ and for any bounded sequences $(u_{n})$ and $(\tilde{u}_{n})$ in $\mathbb{R}^{n}$, the Lipschitz condition
$\sup_{k}\|F_{\Delta t}(u_{n})-F_{\Delta t}(\tilde{u}_{n})\|\leq M\sup_{k}\|u_{n}-\tilde{u}_{n}\|.$ (7)
Note that condition (7) holds if scheme (3) is zero-stable (see e.g. [53]).
In [50], Anguelov and Lubuma observed the elementary stability properties of the linear multistep methods. A numerical method is called elementary stable if the numerical scheme has the same equilibrium points as the continuous model, and their stability is also the same. It was shown that under some mild conditions (with the absolute stability region of the numerical method being bounded, and the original differential equation having at least one stable equilibrium), the ’classical’ linear multistep methods are not elementary stable when their order is greater than one. Because of this, they focused on one-step methods.
In the subsequent section, we demonstrate that under some conditions, the nonstandard linear multistep method (5) attains the same order as the ’standard’ counterpart (3), and it also preserves some key qualitative properties: namely, positivity, a weaker form of monotonicity, and the linear combination of the elements of $u^{n}$.
3 Main results
In this section, we present the main results of the paper. First, the order of the nonstandard multistep methods is observed, and then the preservation of qualitative properties is shown. The proofs of these statements can be found in Section 4.
3.1 The order of nonstandard multistep methods
Before presenting the theorem about the order of nonstandard linear multistep methods, the lemma used in the theorem’s proof is highlighted. The reason for this is that the condition required in the statement of the lemma has substantial implications to be discussed later.
Lemma 1 (Nonstandard form of Taylor’s theorem).
Let $n\geq 0$ be an integer and the functions $f,\varphi:\mathbb{R}\rightarrow\mathbb{R}$ be $n$-many times continuously differentiable on $[a,x]$ and $(n+1)$-many times differentiable on $(a,x)$, where $a,x\in\mathbb{R}$ ($a<x$). Moreover, let us also assume that neither $\varphi$ nor its derivatives vanish on $(a,x)$. Then, the following two claims are equivalent:
  • (C1)
    If $n\geq 1$, then $\varphi(0)=0$, $\varphi^{\prime}(0)=1$, and if $n\geq 2$, then $\varphi^{(k)}(0)=0$ holds for $2\leq k\leq n$.
  • (C2)
    There exists a function $e_{n}(x):\mathbb{R}\rightarrow\mathbb{R}$ such that $\lim_{x\rightarrow a}e_{n}(x)=0$ and
    $f(x)=\sum_{k=0}^{n}\dfrac{f^{(k)}(a)}{k!}\left(\varphi(x-a)\right)^{k}+e_{n}(x)\left(\varphi(x-a)\right)^{n}.$ (8)
    Moreover, $e_{n}(x)=O(|x-a|)$ as $x\rightarrow a$.
Remark 1.
Note that if $\varphi(x)=x$, then we get the ’usual’ form of Taylor’s theorem. Moreover, one could rephrase Lemma 1 in the following way: formula (8) holds if and only if the first $(n+1)$-many terms of the Taylor expansions (around zero) of the functions $\varphi(x)$ and $f(x)=x$ are the same. In Section 3.3 we show several possible choices of $\varphi$ for which condition (C1) holds for different values of $n$, and these choices are compared in Section 5.
Remark 2.
It is worth mentioning here that one can generalize the previous lemma in the following way: let us assume that (other than the smoothness assumption in the statement of Lemma 1) $\varphi(0)=0$, $\varphi(x)>0$ as $x>0$, and $\varphi^{\prime}(x)\neq 0$ as $x>0$. Then, a similar form can be proven, namely
$f(x)$ $=\sum_{k=0}^{n}\dfrac{f^{(k)}(a)}{k!}\left(\varphi(x-a)\right)^{k}+\dfrac{f^{(n+1)}(\xi)}{(n+1)!}\dfrac{\left(\varphi(x-a)\right)^{n+1}}{\varphi^{\prime}(x-\xi)}+$ (9)
$+\sum_{k=0}^{n-1}\dfrac{f^{(k+1)}(\xi)}{k!}\left(\varphi(x-\xi)\right)^{k}\left(1-\varphi^{\prime}(x-\xi)\right),$
where $\xi\in(a,x)$. (Note that for $\varphi(x)=x$, this is Taylor’s theorem with the Lagrange form of the remainder.) Nevertheless, because of the last term of (9), this form is not beneficial in our case.
Now, we state the main result about the order of nonstandard multistep methods.
Theorem 2.
Let us assume that for equation (1) with a smooth solution $y\in C^{p+1}([t_{0},T])$, the ’standard’ linear multistep method (3) is convergent of order $p$. Let us assume that for function $\varphi(x)$, it is $(p+1)$-times continuously differentiable, and neither $\varphi$ nor its derivatives vanish for $x>0$. Then, the nonstandard linear multistep method (5) is convergent of order $p$ if and only if condition (C1) of Lemma 1 holds with $n=p$.
3.2 Preservation of qualitative properties
In this section, we observe the preservation of some qualitative properties of the original differential equation (1). In all three cases, we consider the Forward Euler method applied to equation (1) with timestep $\Delta t_{FE}$, and denote the resulting numerical scheme by $v^{n}=(v_{1}^{n},\;v_{2}^{n},\;\dots,\;v_{m}^{n})^{T}$.
Then, we assume that this method preserves some qualitative property under some conditions for the timestep, and then observe the same property for the explicit nonstandard multistep method with $s$-many steps written in the form
$u^{n+1}=\sum_{j=1}^{s}\left(\tilde{\alpha}_{j}u^{n+1-j}+\varphi(\Delta t)\tilde{\beta}_{j}f(u^{n+1-j})\right),$ (10)
where we assume that $\tilde{\alpha}_{j},\tilde{\beta}_{j}\geq 0$ and if $\tilde{\alpha}_{j}=0$, then $\tilde{\beta}_{j}=0$. Both the Euler method and the multistep methods are started from $v^{0}=u^{0}=\tilde{y}=y(0)$.
First, we focus on the boundedness of solutions. We say that a component $y_{j}(t)$ of the solution $y(t)=(y_{1}(t),\;y_{2}(t),\;\dots,\;y_{m}(t))^{T}$ is bounded from above, if $y_{j}(t)\leq M$ holds for every $t\in[t_{0},T]$ for some $M\in\mathbb{R}$. The boundedness from below is also similar.
[…]
Let us assume that for component $y_{k}(t)$ of the solution of equation (1) we have $y_{k}(t)\leq M$ for $t\in[t_{0},T]$. The next theorem states that in this case (under some conditions) the numerical method (10) also preserves this property for any choice of the step size.
Theorem 3.
Let us assume that the following conditions are met:
  • If $\Delta t_{FE}\leq B_{FE}$ for some constant $B_{FE}\in\mathbb{R}^{+}$, then the Forward Euler method applied to (1) preserves the boundedness property, namely $v_{k}^{1}\leq M$ holds.
  • For the starting values of the multistep method (10), we have $u_{k}^{0},u_{k}^{1},\dots,u_{k}^{s-1}\leq M$.
  • $\varphi(x)\leq\mathcal{C}B_{FE}$, where $\mathcal{C}\in\mathbb{R}^{+}$ is the SSP coefficient defined as (see [52])
    $\mathcal{C}=\min_{j}\dfrac{\tilde{\alpha}_{j}}{\tilde{\beta}_{j}}.$
Then, for the nonstandard multistep method (10), $u_{k}^{n}\leq M$ also holds for $n=s,\dots\mathcal{N}$ for any stepsize $\Delta t\geq 0$.
The same theorem also holds if we consider boundedness from below.
Another theorem can be stated if we consider the monotone property of the solution. Let us assume that the component $y_{k}(t)$ of the solution of equation (1) is monotone increasing for $t\in[t_{0},T]$ if $\tilde{y}\in(A,B)$ element-wise ($A,B\in\mathbb{R},\;A<B$, where we also allow that $A=-\infty$ or $B=\infty$).
Theorem 4.
Let us assume that the following conditions are met:
  • If $\Delta t_{FE}\leq B_{FE}$ for some constant $B_{FE}\in\mathbb{R}^{+}$, then the Forward Euler method applied to (1) preserves the monotone increase property, namely $v_{k}^{1}\geq v_{k}^{0}$, and $v_{k}^{1}\in(A,B)$ holds if $v_{k}^{0}\in(A,B)$.
  • For the starting values of the multistep method (10), we have $u_{k}^{0},\;u_{k}^{1},\;\dots,\;u_{k}^{s-1}\in(A,B)$.
  • $\varphi(x)\leq\mathcal{C}B_{FE}$, where $\mathcal{C}\in\mathbb{R}^{+}$ is the SSP coefficient defined as in Theorem 3.
Then, for the nonstandard multistep method (10),
$\min_{\ell=n-s+1,\;n-s+2,\;\dots,\;n}\{u_{k}^{\ell}\}\leq u_{k}^{n+1}$ (11)
also holds for $n=s,\dots\mathcal{N}$ for any stepsize $\Delta t\geq 0$.
An analogous theorem can be proven if we consider monotone decrease: namely, the final inequality in this case takes the form
$\max_{\ell=n-s+1,\;n-s+2,\;\dots,\;n}\{u_{k}^{\ell}\}\geq u_{k}^{n+1}.$
Remark 3.
Note that property (11) is only a necessary, but not sufficient condition for the monotone increase of numerical solution $u_{k}^{n}$. Numerical experiments indicate that the numerical scheme can indeed preserve the monotone property if the bound for $\varphi(x)$ is even smaller than $\mathcal{C}B_{FE}$, but for $\varphi(x)\leq\mathcal{C}B_{FE}$, only inequality (11) holds (see the tests in Section 5). The determination of a bound under which the monotonicity is also preserved could be a topic of further research.
The next theorem is about a property concerning the preservation of some linear combination of the elements of $u^{n}$. Let us assume that if $y(t_{0})\in S$ for some $S\subset\mathbb{R}^{m}$ hyperrectangle, then for some linear function $P:\mathbb{R}^{m}\rightarrow\mathbb{R}$ we have
$P(y_{1}(t),\;y_{2}(t),\;\dots,y_{m}(t))=\mathcal{M}\qquad\forall t\in[t_{0},\;T]$
for some $\mathcal{M}\in\mathbb{R}$.
Theorem 5.
Let us assume that the following conditions are met:
  • Consider the Forward Euler method starting from $v^{0}\in S$. If $\Delta t_{FE}\leq B_{FE}$ for some constant $B_{FE}\in\mathbb{R}^{+}$, then $P(v_{1}^{n},\;v_{2}^{n},\;\dots,v_{m}^{n})=\mathcal{M}$ for $n=1,2,\dots$ assuming that $P(v_{1}^{0},\;v_{2}^{0},\;\dots,v_{m}^{0})=\mathcal{M}$.
[…]
  • $\varphi(x)\leq\mathcal{C}B_{FE}$, where $\mathcal{C}\in\mathbb{R}^{+}$ is the SSP coefficient defined as in Theorem 3.
Then,
$P(v_{1}^{n},\;v_{2}^{n},\;\dots,v_{m}^{n})=\mathcal{M}$
also holds for $n=s,\dots\mathcal{N}$.
Remark 4.
If $P(y_{1}(t),\;y_{2}(t),\;\dots,\;y_{m}(t))=\sum_{j=1}^{m}y_{j}(t)$, then the previous theorem is about the preservation of the sum of all components (see Section 5.2 for such an example).
3.3 The choice of function $\varphi$
Here, we briefly address the issue of the choice of function $\varphi$. As it was mentioned in the previous sections, for a nonstandard multistep method to be of order $p$ and to preserve some important qualitative properties, the following two conditions should be met:
  • (A)
    $\varphi$ should be $(p+1)$-times continuously differentiable (for $x>0$) and $\varphi(x)>0$ for $x>0$. Moreover, if $p\geq 1$, then $\varphi(0)=0$, $\varphi^{\prime}(0)=1$ and if $p\geq 2$ then $\varphi^{(k)}(0)=0$ holds for $2\leq k\leq p$.
  • (B)
    $\varphi(x)\leq\mathcal{B}:=\mathcal{C}B_{FE}$ for every $x>0$.
In the following, we outline different possible choices for $\varphi$. These are chosen with condition (B) fulfilled, and they are arranged according to the value $p$ for which condition (A) is satisfied.
  • With $p=1$:
    • $\varphi_{1}(x)=\mathcal{B}(1-e^{-\frac{x}{\mathcal{B}}})$ (here $\varphi^{\prime\prime}(0)=-\frac{1}{\mathcal{B}}\neq 0$),
    • $\varphi_{2}(x)=xe^{-\frac{x}{\mathcal{B}e}}$ (here $\varphi^{\prime\prime}(0)=-\frac{2}{\mathcal{B}e}\neq 0$),
    • $\varphi_{3}(x)=\dfrac{\mathcal{B}x}{\mathcal{B}+x}$ (here $\varphi^{\prime\prime}(0)=-\frac{2}{\mathcal{B}}\neq 0$).
  • With $p=2$:
    • $\varphi_{4}(x)=\dfrac{\mathcal{B}2}{\pi}\mathrm{arctan}\left(\dfrac{x\pi}{2\mathcal{B}}\right)$ (here $\varphi^{\prime\prime\prime}(0)=-\dfrac{\pi^{2}}{2\mathcal{B}^{2}}\neq 0$),
    • $\varphi_{5}(x)=\mathcal{B}\mathrm{tanh}\left(\dfrac{x}{\mathcal{B}}\right)$ (here $\varphi^{\prime\prime\prime}(0)=-\dfrac{2}{\mathcal{B}^{2}}\neq 0$).
    • $\varphi_{6}(x)=\dfrac{\mathcal{B}x}{\left(\mathcal{B}^{2}+x^{2}\right)^{1/2}}$ (here $\varphi^{\prime\prime\prime}(0)=-\frac{3}{\mathcal{B}^{2}}\neq 0$).
  • With $p=3$: $\varphi_{7}(x)=\dfrac{\mathcal{B}x}{\left(\mathcal{B}^{3}+x^{3}\right)^{1/3}}$ (here $\varphi^{(4)}(0)=-\dfrac{8}{\mathcal{B}^{3}}\neq 0$).
  • With $p=4$: $\varphi_{8}(x)=\dfrac{\mathcal{B}x}{\left(\mathcal{B}^{4}+x^{4}\right)^{1/4}}$ (here $\varphi^{(5)}(0)=-\dfrac{30}{\mathcal{B}^{4}}\neq 0$).
  • For the general $p\geq 5$ case: $\varphi(x)=\dfrac{\mathcal{B}x}{\left(\mathcal{B}^{p}+x^{p}\right)^{1/p}}$ (here $\varphi^{(p+1)}(0)=-\dfrac{(p-1)!+p!}{\mathcal{B}^{p}}\neq 0$).
In Section 5 we perform several numerical experiments comparing these choices. We should also note here that in [49], Dang and Hoang demonstrated a different construction of functions fulfilling these conditions up to an arbitrary order.
4 Proofs of the statements of Section 3
In this section, we present the proofs of the statements of the previous section.
Proof of Lemma 1.
We prove the two directions of the statement separately.
  1. 1.
    Let us start with the (C1$\Rightarrow$(C2) direction.
    We prove the statement by induction: first, we show that it holds for $n=0$ and $n=1$ (the case of $n=2$ is also presented for easier understanding), and then by assuming that it holds for $n=N$, we validate the case $n=N+1$.
    If $n=0$, (8) has the form $f(x)=f(a)+e_{0}(x)$, meaning that $\lim_{x\rightarrow a}e_{0}(x)=\lim_{x\rightarrow a}f(x)-f(a)=0$.
    If $n=1$, (8) has the form $f(x)=f(a)+f^{\prime}(a)\varphi(x-a)+e_{1}(x)\varphi(x-a)$, from which we get
    $\lim_{x\rightarrow a}e_{1}(x)=\lim_{x\rightarrow a}\dfrac{f(x)-\left(f(a)+f^{\prime}(a)\varphi(x-a)\right)}{\varphi(x-a)}=$
    Now we can use the rule of l’Hospital, since $\varphi(0)=0$:
    $=\lim_{x\rightarrow a}\dfrac{f^{\prime}(x)-f^{\prime}(a)\varphi^{\prime}(x-a)}{\varphi^{\prime}(x-a)}=0,$
    where we used that $\varphi^{\prime}(0)=1$.
    If $n=2$, (8) has the form
    $f(x)=f(a)+f^{\prime}(a)\varphi(x-a)+\dfrac{f^{\prime\prime}(a)}{2}\left(\varphi(x-a)\right)^{2}+e_{2}(x)\left(\varphi(x-a)\right)^{2},$
    from which we get
    $\lim_{x\rightarrow a}e_{2}(x)=\lim_{x\rightarrow a}\dfrac{f(x)-\left(f(a)+f^{\prime}(a)\varphi(x-a)+\frac{f^{\prime\prime}(a)}{2}\left(\varphi(x-a)\right)^{2}\right)}{\left(\varphi(x-a)\right)^{2}}=$
    Now we can use the rule of l’Hospital, since $\varphi(0)=0$:
    $=\lim_{x\rightarrow a}\dfrac{f^{\prime}(x)-f^{\prime}(a)\varphi^{\prime}(x-a)-f^{\prime\prime}(a)\varphi(x-a)\varphi^{\prime}(x-a)}{2\varphi(x-a)\varphi^{\prime}(x-a)}=$
    Here we use l’Hospital’s rule again:
    $=\lim_{x\rightarrow a}\dfrac{f^{\prime\prime}(x)-f^{\prime}(a)\varphi^{\prime\prime}(x-a)-f^{\prime\prime}(a)\left(\left(\varphi^{\prime}(x-a)\right)^{2}+\varphi(x-a)\varphi^{\prime\prime}(x-a)\right)}{2\left(\varphi^{\prime}(x-a)\right)^{2}+2\varphi(x-a)\varphi^{\prime\prime}(x-a)}=0,$
    where we used that $\varphi(0)=0$, $\varphi^{\prime}(0)=1$ and $\varphi^{\prime\prime}(0)=0$.
    Now let us assume that (C2) holds for $n=N$. Observe that it is enough to show that
    $\lim_{x\rightarrow a}e_{N+1}(x)=-\dfrac{f^{\prime}(a)\varphi^{(N+1)}(0)}{(N+1)!}.$ (12)
    From this, if $\varphi^{(N+1)}(0)=0$ holds, then the statement is proven.
    Let $k<N+1$ be a positive integer and observe the limit $\lim_{x\rightarrow a}e_{N+1}$ after $k$-many applications of the l’Hospital rule. Note that apart from the terms containing $f^{(N+1)}(a)$, the numerator is the same as the numerator of $e_{N}$ after the same number of applications of the rule. This means that by the assumption of the induction step, these terms tend to zero. Moreover, the terms containing $f^{(N+1)}(a)$ initially had the ($N+1$)th power of $\varphi(x-a)$, meaning that after $k$-many ($k<N+1$) differentiations every term will contain at least one $\varphi(x-a)$ term, meaning that this also tends to zero (since $\varphi(0)=0$). The same can also be applied to the denominator, meaning that (at least) one additional l’Hospital rule can be applied. This means that in order to calculate $\lim_{x\rightarrow a}e_{N+1}$, $N+1$-many l’Hospital rules have to be applied.
    Now, observe the form of the limit after ($N+1$)-many differentiations. In the numerator, almost every term has at least either one $\varphi(x-a)$ term or one $\varphi^{(\ell)}(x-a)$ ($2\leq\ell\leq N$) term, thus, by the assumptions of (C1), all of these tend to zero. The three exceptions are:
    • the term $f^{(N+1)}(x)$,
    • the term $-f^{\prime}(a)\varphi^{(N+1)}(x-a)$,
    • the term $\left(\varphi^{\prime}(x-a)\right)^{N+1}$. (The case $\left(\varphi^{\prime}(x-a)\right)^{\ell}$ where $\ell<N+1$ is not possible since we performed $N+1$-many differentiations, so the sum of the orders of derivatives in a product should also be $N+1$.) Such a term is multiplied by $-f^{(N+1)}(a)$ (it cannot appear elsewhere since the number of terms in a product remains the same during differentiation).
    Note that the denominator is similar to the terms in the numerator multiplied by $f^{(N+1)}(a)$: in the denominator, instead of $f^{(N+1)}(a)$, they are multiplied by $(N+1)!$. Consequently, a similar argument can also be made there: the only nonzero term is $\left(\varphi^{\prime}(x-a)\right)^{N+1}$.
    In conclusion,
    $\lim_{x\rightarrow a}e_{N+1}(x)=\dfrac{f^{(N+1)}(a)-f^{\prime}(a)\varphi^{(N+1)}(0)-f^{(N+1)}(a)\left(\varphi^{\prime}(0)\right)^{N+1}}{(N+1)!\left(\varphi^{\prime}(0)\right)^{N+1}}=-\dfrac{f^{\prime}(a)\varphi^{(N+1)}(0)}{(N+1)!}.$
    Therefore, claim (12) is proven.
    Now we show that $e_{n}(x)=O(|x-a|)$. For this, it is enough to prove that $\dfrac{|e_{n}(x)|}{|x-a|}$ is bounded if $|x-a|$ is smaller than a given constant. Let us observe that
    $\dfrac{|e_{n}(x)|}{|x-a|}=\dfrac{\left|f(x)-f(a)-f^{\prime}(a)\varphi(x-a)-\dfrac{f^{\prime\prime}(a)}{2}\left(\varphi(x-a)\right)^{2}-\dots\dfrac{f^{(n)}(a)}{n!}\left(\varphi(x-a)\right)^{n}\right|}{|x-a|}$ (13)
    By condition (C1), we get that the ’standard’ Taylor series of $\varphi(x-a)$ around 00 has the form $\varphi(x-a)=(x-a)+O(|x-a|^{n+1})$, meaning that the numerator of (13) has the form
    $\left|f(x)-f(a)-f^{\prime}(a)\left((x-a)+O(|x-a|^{n+1})\right)-\dfrac{f^{\prime\prime}(a)}{2}\left((x-a)+O(|x-a|^{n+1})\right)^{2}-\dots\right.$
    $\left.\dots-\dfrac{f^{(n)}(a)}{n!}\left((x-a)+O(|x-a|^{n+1})\right)^{n}\right|=$
    $=\left|\dfrac{f^{(n+1)}(\xi_{1})}{(n+1)!}|x-a|^{n+1}-f^{\prime}(a)\dfrac{\varphi^{(n+1)}(\xi_{2})}{(n+1)!}|x-a|^{n+1}+O(|x-a|^{n+2})\right|,$
    where $\xi_{1},\xi_{2}\in(a,x)$ and we used the Lagrange forms of the remainders. Then, by dividing this term by $|x-a|$, we get an expression that is bounded if $|x-a|$ is small enough.
  2. 2.
    Here we show the (C2$\Rightarrow$(C1) direction.
    We again proceed by induction. For $n=0$, we have no condition.
    For $n=1$, we assume proceeding towards contradiction that $\varphi(0)\neq 0$. Then, (8) has the form
    $f(x)=f(a)+f^{\prime}(a)\varphi(x-a)+e_{1}(x)\varphi(x-a).$
    By taking the limit $x\rightarrow a$ on both sides, we get a contradiction.
    Now, assume that $\varphi^{\prime}(0)\neq 1$. Similarly to the proof of the first part of the statement, we get that
    $0=\lim_{x\rightarrow a}e_{1}(x)=\dfrac{f^{\prime}(a)-f^{\prime}(a)\varphi^{\prime}(0)}{\varphi^{\prime}(0)}=f^{\prime}(a)\left(\dfrac{1}{\varphi^{\prime}(0)}-1\right).$
    This can only hold for any arbitrary $f$ (which has the properties specified in the statement of the lemma) if $\varphi^{\prime}(0)=1$.
    In order to prove the induction step, let us assume that the statement is true for $n=N$. As in the previous half of the proof, one can show that
    $\lim_{x\rightarrow a}e_{N+1}(x)=-\dfrac{f^{\prime}(a)\varphi^{(N+1)}(0)}{(N+1)!},$
    which again can only be zero if $\varphi^{(N+1)}(0)=0$. ∎
Remark 5.
Claim (12) can also be proven by applying the Faà di Bruno formula (see e.g. [54]) to $g(h(x))$ with $g(x)=x^{N+1}$ and $h(x)=\varphi(x-a)$. Moreover, the direction (C1) $\Rightarrow$ (C2) is also the consequence of the proof of the $e_{n}(x-a)=O(|x-a|)$ claim, but the longer version is needed for the (C2) $\Rightarrow$ (C1) direction.
Proof of Theorem 2.
We prove the two directions separately.
  1. 1.
    First we show that from condition (C1), we get a convergence of order $p$.
    Let us observe the following linear difference operator $L$ (see e.g. [51]):
    $L(y,t,\Delta t)=\sum_{j=0}^{k}\left(\alpha_{j}y(t+j\Delta t)-\varphi(\Delta t)\beta_{j}y^{\prime}(t+j\Delta t)\right),$
    which we got by substituting the exact solution $y(t)$ into scheme (5). Now, let us apply the nonstandard Taylor expansion (around $a=0$) defined in Lemma 1 for functions $y$ and $y^{\prime}$ at points $x=t+j\Delta t$:
    $L(y,t,\Delta t)=\sum_{j=0}^{k}\alpha_{j}\sum_{q=0}^{p}\dfrac{\left(\varphi(j\Delta t)\right)^{q}}{q!}y^{(q)}(t)-\varphi(\Delta t)\sum_{j=0}^{k}\beta_{j}\sum_{r=0}^{p}\dfrac{\left(\varphi(j\Delta t)\right)^{r}}{r!}y^{(r+1)}(t)+O((\Delta t)^{p+1})=$
    $=y(t)\sum_{j=0}^{k}\alpha_{j}+\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\sum_{j=0}^{k}\alpha_{j}\left(\varphi(j\Delta t)\right)^{q}-\varphi(\Delta t)\sum_{r=0}^{p}\dfrac{y^{(r+1)}(t)}{r!}\sum_{j=0}^{k}\beta_{j}\left(\varphi(j\Delta t)\right)^{r}+O((\Delta t)^{p+1})=$
    $=\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\sum_{j=0}^{k}\alpha_{j}\left(\varphi(j\Delta t)\right)^{q}-\varphi(\Delta t)\sum_{r=1}^{p+1}r\dfrac{y^{(r)}(t)}{r!}\sum_{j=0}^{k}\beta_{j}\left(\varphi(j\Delta t)\right)^{r-1}+O((\Delta t)^{p+1}),$
    where we used that the scheme (3) is consistent, meaning that $\sum_{j=0}^{k}\alpha_{j}=0$ holds. Moreover, if condition (C1) is true for $n=p$, then it means that $\varphi(h)=h+O(h^{p+1})$, which, substituted into the previous formula, gives
    $\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\left(\Delta t\right)^{q}\sum_{j=0}^{k}\alpha_{j}j^{q}-\sum_{r=1}^{p+1}r\dfrac{y^{(r)}(t)}{r!}\left(\Delta t\right)^{r}\sum_{j=0}^{k}\beta_{j}j^{r-1}+O((\Delta t)^{p+1})=$
    $=\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\left(\Delta t\right)^{q}\left(\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}\right)+O((\Delta t)^{p+1}).$
    If scheme (3) has an order of $p$, then $\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}=0$ for $q=1,\dots p$ (see e.g. [51], Theorem 2.4.), meaning that the nonstandard method also has an order of $p$.
  2. 2.
    Now we show that an order of $p$ can only hold if (C1) is satisfied with $n=p$.
    We prove the statement by induction. Let us consider the case $n=1$, i.e., the method’s order is at least one. It is obvious that consistency requires $\varphi(0)=0$. Moreover, if we consider the Taylor expansion of $\varphi$ around zero, then (since $\varphi(0)=0$) we have $\varphi(\Delta t)=\varphi^{\prime}(0)\Delta t+O((\Delta t)^{2})$. By substituting this into operator $L$ (defined in the first part of the proof), we get
    $L(y,t,\Delta t)=\sum_{j=0}^{k}\left(\alpha_{j}y(t+j\Delta t)-\varphi^{\prime}(0)\Delta t\beta_{j}y^{\prime}(t+j\Delta t)\right)+O((\Delta t)^{2}).$
    Let us apply the ’standard’ Taylor expansion for the functions $y(t+j\Delta t)$ and $y^{\prime}(t+j\Delta t)$ at points $x=t+j\Delta t$:
    $L(y,t,\Delta t)=\sum_{j=0}^{k}\alpha_{j}\left(y(t)+j\Delta ty^{\prime}(t)\right)-\varphi^{\prime}(0)\Delta t\sum_{j=0}^{k}\beta_{j}\left(y^{\prime}(t)+j\Delta ty^{\prime\prime}(t)\right)+O((\Delta t)^{2})=$
    $=\sum_{j=0}^{k}\alpha_{j}y(t)+y^{\prime}(t)\Delta t\sum_{j=0}^{k}\left(j\alpha_{j}-\varphi^{\prime}(0)\beta_{j}\right)+O((\Delta t)^{2})=$
    $=y^{\prime}(t)\Delta t\sum_{j=0}^{k}\left(j\alpha_{j}-\varphi^{\prime}(0)\beta_{j}\right)+O((\Delta t)^{2}),$
    where we used that $\sum_{j=0}^{k}\alpha_{j}=0$ holds. Moreover, we also know that $\sum_{j=0}^{k}\alpha_{j}j-\sum_{j=0}^{k}\beta_{j}=0$, meaning that the above expression equals $O((\Delta t)^{2})$ only if $\varphi^{\prime}(0)=1$.
    Now, let us assume that condition (C1) holds for $n<N$ and we prove the $n=N$ case ($2\leq N\leq p$). The argument is rather similar to the one in the $n=1$ step: in this case $\varphi(\Delta t)=\Delta t+\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}+O((\Delta t)^{N+1})$. By substitution, we get
    $L(y,t,\Delta t)=\sum_{j=0}^{k}\left(\alpha_{j}y(t+j\Delta t)-\left(\Delta t+\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\right)\beta_{j}y^{\prime}(t+j\Delta t)\right)+O((\Delta t)^{N+1}).$
    As before, we apply the ’standard’ Taylor expansion to functions $y(t+j\Delta t)$ and $y^{\prime}(t+j\Delta t)$ at points $x=t+j\Delta t$, resulting in
    $L(y,t,\Delta t)$ $=\sum_{j=0}^{k}\left(\alpha_{j}\sum_{q=0}^{N}\dfrac{\left(j\Delta t\right)^{q}}{q!}y^{(q)}(t)-\right.$
    $-\left.\left(\Delta t+\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\right)\beta_{j}\sum_{r=0}^{N}\dfrac{\left(j\Delta t\right)^{r}}{r!}y^{(r+1)}(t)\right)+O((\Delta t)^{N+1})=$
    $=\sum_{q=1}^{N}\dfrac{y^{(q)}(t)}{q!}\left(\Delta t\right)^{q}\left(\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}\right)-\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\sum_{j=0}^{k}\beta_{j}y^{\prime}(t)+O((\Delta t)^{N+1})=$
    $=-\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\sum_{j=0}^{k}\beta_{j}y^{\prime}(t)+O((\Delta t)^{N+1}),$ (14)
    where we used that $\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}=0$ for $q=1,\dots N$ ($N\leq p$) and $\sum_{j=0}^{k}\alpha_{j}=0$. Therefore, expression (14) can only equal $O((\Delta t)^{N+1})$ in general if $\varphi^{(N)}(0)=0$. This concludes this part of the proof. ∎
Proof of Theorem 3.
We prove the statement by induction: let us assume that $u_{k}^{0},\;u_{k}^{1},\;\dots u_{k}^{N}\leq M$ and we aim to show that $u_{k}^{N+1}\leq M$ (here $N+1\leq\mathcal{N}$).
Let us consider the numerical method (10) and observe its part concerning the element $u_{k}$:
$u_{k}^{N+1}=\sum_{j=1}^{s}\tilde{\alpha}_{j}\left(u_{k}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{k}(u^{N+1-j})\right),$ (15)
where we used the notation $f_{k}$ for the $k$th element of the function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$. Since the boundedness property holds for the forward Euler method if $\Delta t_{FE}\leq B_{FE}$, for $\Delta t_{FE}\leq B_{FE}$ we have
$\left(v_{k}^{0}+\Delta tf_{k}(v^{0})\right)\leq M,$
where $v_{k}^{0}$ is an arbitrary value for which $v_{k}^{0}\leq M$ holds. By the assumption of the theorem, $\varphi(\Delta t)\leq\mathcal{C}B_{FE}$ is true, meaning that for every $j$
$\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}\leq\mathcal{C}B_{FE}\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}\leq B_{FE}.$
Therefore, every term of the sum (15) can be thought of as a Forward Euler step with stepsize $\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}\leq B_{FE}$, resulting in
$\sum_{j=1}^{s}\tilde{\alpha}_{j}\left(u_{k}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{k}(u^{N+1-j})\right)\leq\sum_{j=1}^{s}\tilde{\alpha}_{j}M=M,$
where we used that $\sum_{j=1}^{s}\tilde{\alpha}_{j}=1$ (this holds because the method is consistent, see [52]). ∎
Proof of Theorem 4.
We prove the statement by induction: let us assume that $u_{k}^{n+1}\geq\min_{\ell=n-s+1,\;n-s+2,\;\dots,\;n}\{u_{k}^{\ell}\}$ for $n=s-1,\;s,\;\dots,\;N-1$ and our goal is to show that $u_{k}^{N+1}\geq\min_{\ell=N-s+1,\;N-s+2,\;\dots,\;N}\{u_{k}^{\ell}\}$ ($N+1\leq\mathcal{N}$).
We again consider the numerical method (10) and observe its part concerning the element $u_{k}$ (see (15) in the proof of Theorem 3). Since the monotone increase holds for the forward Euler method if $\Delta t_{FE}\leq B_{FE}$, for $\Delta t_{FE}\leq B_{FE}$ we have
$v_{k}^{0}+\Delta tf_{k}(v^{0})\geq v_{k}^{0},$
where $v_{k}^{0}$ is an arbitrary value for which $v_{k}^{0}\in(A,B)$ holds. By the assumption of the theorem, $\varphi(\Delta t)\leq\mathcal{C}B_{FE}$ is true, meaning that for every $j$, $\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}\leq B_{FE}$. Moreover, by the statement of Theorem 3, $u_{k}^{n}\in(A,B)$ for $n=s-1,\;s,\;\dots,\;N-1$. Therefore, every term of the sum (15) can be thought of as a Forward Euler step with stepsize $\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}\leq B_{FE}$, resulting in
$u_{k}^{N+1}=\sum_{j=1}^{s}\tilde{\alpha}_{j}\left(u_{k}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{k}(u^{N+1-j})\right)\geq\sum_{j=1}^{s}\tilde{\alpha}_{j}u_{k}^{N+1-j}.\geq$
$\geq\sum_{j=1}^{s}\tilde{\alpha}_{j}\min_{\ell=N+1-s,\;N+2-s,\;\dots,\;N}\{u_{k}^{\ell}\}\geq\min_{\ell=N-s+1,\;N-s+2,\;\dots,\;N}\{u_{k}^{\ell}\},$
where we used that $\sum_{j=1}^{s}\tilde{\alpha}_{j}=1$. ∎
Proof of Theorem 5.
We prove the statement by induction: let us assume that $P(u_{1}^{n},\;u_{2}^{n},\;\dots,u_{m}^{n})=\mathcal{M}$ for $n=s,\;s+1,\;\dots,\;N$ and our goal is to show that $P(u_{1}^{N+1},\;u_{2}^{N+1},\;\dots,u_{m}^{N+1})=\mathcal{M}$ ($N+1\leq\mathcal{N}$).
Let us consider the numerical method (10) and observe its part concerning the element $u_{k}$ (see (15) in the proof of Theorem 3). Since the preservation of the value $\mathcal{M}$ holds for the forward Euler method if $\Delta t_{FE}\leq B_{FE}$, for $\Delta t_{FE}\leq B_{FE}$ we have
$P(v_{1}^{N+1},\;v_{2}^{N+1},\;\dots,v_{m}^{N+1})=\sum_{\ell=1}^{m}\gamma_{\ell}v_{\ell}^{N+1}=\sum_{\ell=1}^{m}\gamma_{\ell}\left(v_{\ell}^{N}+\Delta tf_{\ell}(v^{N})\right)=\sum_{\ell=1}^{m}\gamma_{\ell}v_{\ell}^{N}=M,$
where $\gamma_{\ell}\in\mathbb{R}$ are some given constants.
By the assumption of the theorem, $\varphi(\Delta t)\leq\mathcal{C}B_{FE}$ is true, meaning that for every $j$, $\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}\leq B_{FE}$. Therefore, every term of the sum (15) can be thought of as a Forward Euler step with stepsize $\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}\leq B_{FE}$, resulting in
$P(u_{1}^{N+1},\;u_{2}^{N+1},\;\dots,u_{m}^{N+1})=\sum_{\ell=1}^{m}\gamma_{\ell}u_{\ell}^{N+1}=\sum_{\ell=1}^{m}\gamma_{\ell}\sum_{j=1}^{s}\tilde{\alpha}_{j}\left(u_{\ell}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{\ell}(u^{N+1-j})\right)=$
$=\sum_{j=1}^{s}\tilde{\alpha}_{j}\sum_{\ell=1}^{m}\gamma_{\ell}\left(u_{\ell}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{\ell}(u^{N+1-j})\right)=\sum_{j=1}^{s}\tilde{\alpha}_{j}M=M,$
where we used that $\sum_{j=1}^{s}\tilde{\alpha}_{j}=1$. ∎
5 Applications and numerical simulations
In this section, we apply the previously observed nonstandard multistep method (10) to different initial value problems.
Prior to showing some specific applications, the main steps of the construction of property-preserving nonstandard multistep methods are outlined.
  1. 1.
    Assume that the original differential equation (1) attains one of the qualitative properties mentioned in Section 3.2, which should be preserved by the numerical method too.
  2. 2.
    Apply the Forward Euler method to the given equation with $\Delta t$ and calculate such a bound $B_{FE}$, for which the observed property is preserved if $\Delta t\leq B_{FE}$ holds.
  3. 3.
    Choose function $\varphi(x)$ in a way that $\varphi(x)\leq\mathcal{C}B_{FE}$. For this, one of the possible choices listed in Section 3.3 can be used with $\mathcal{B}=\mathcal{C}B_{FE}$.
  4. 4.
    For a method with $s$-many steps, the values $u^{1},\;u^{2},\;\dots,\;u^{s-1}$ should be calculated. For this, a nonstandard Runge-Kutta method can be used with an appropriate function $\varphi(x)$ (see example in Section 5.2 for details).
  5. 5.
    Thus, the nonstandard multistep method in the form (10) preserves the desired property and has the same order as the standard counterpart.
In Section 5.1, the method is applied to a one-dimensional problem, while in Section 5.2 a system is considered.
5.1 One-dimensional logistic growth equation
Consider the one-dimensional ($n=1$) logistic growth equation
$\begin{cases}y^{\prime}(t)&=y(t)\left(2-y(t)\right),\\ y(0)&=\tilde{y}.\end{cases}$ (16)
The solution of the above equation is $y(t)=\dfrac{2\tilde{y}}{\tilde{y}+(2-\tilde{y})e^{-2x}}$. It is also apparent that the two equilibria of the equation are $y_{1}^{*}=0$ and $y_{2}^{*}=2$, where $y_{1}^{*}$ is unstable and $y_{2}^{*}$ is asymptotically stable. The solution of the equation is also monotone.
To apply the nonstandard multistep method, we should first observe the behavior of the Forward Euler method.
Proposition 1.
The Forward Euler method applied to equation (16) preserves the monotonicity and the boundedness of the solution if $\Delta t\leq\min\left\{\dfrac{1}{2},\dfrac{1}{\tilde{y}}\right\}$ if $\tilde{y}\geq 0$, and if $\tilde{y}<0$, then the method preserves the properties unconditionally.
Proof.
Let us apply the forward Euler method to equation (16):
$u^{n+1}=u^{n}+\Delta tu^{n}(2-u^{n})=u^{n}(1+2\Delta t-\Delta t\;u^{n}).$
We prove the statement by induction: let us assume that the method behaves as expected for $n=0,\dots,N$. Consider four different cases:
  • If either $\tilde{y}=0$ or $\tilde{y}=2$, the numerical solution remains constant.
  • Let us assume that $\tilde{y}=u^{0}<0$. In this case, the solution should remain negative, and it should be monotone decreasing. If $u^{n}<0$, then $u^{n+1}$ is also negative if $01+2\Delta t-\Delta t\;u^{n}>$ 0, but this is true. For the decreasing property, we need that $11+2\Delta t-\Delta t\;u^{n}>$ 1, but this holds too. Consequently, in this case, the method preserves the boundedness and the monotonicity of the solution unconditionally.
  • Assume that $20<\tilde{y}<$ 2. For the increasing property, it is enough to show that $11+2\Delta t-\Delta t\;u^{n}>$ 1, but this is true since $u^{n}>0$. For the boundedness from above, consider the function $g(x)=x(1+2\Delta t-\Delta t\;x)$. This attains its maximum at $\dfrac{1+2\Delta t}{2\Delta t}$, and its value at this point is $\dfrac{(1+2\Delta t)^{2}}{4\Delta t}$ which is bigger than $2$ for every positive value of $\Delta t$. Because of this, let us choose $\Delta t$ in a way that the maximum point is bigger than $2$: $\Delta t<\frac{1}{2}$ is sufficient. Therefore, the function $g$ is monotone increasing on $[0,2]$, with its maximum point at $x=2$ with value $g(2)=2$. Therefore, $u^{n+1}=g(u^{n})<2$ if $\Delta t<\dfrac{1}{2}$. The boundedness from below is trivial because of the monotone increasing property.
  • Consider the case $\tilde{y}>2$. For the decreasing property we need $11+2\Delta t-\Delta t\;u^{n}<$ 1 but this is true since $u^{n}>2$. For the boundedness from below, consider the function $g(x)=x(1+2\Delta t-\Delta t\;x)$. We should prove that if $\tilde{y}>x>2$, then $g(x)>2$. In the previous case we have shown that if $\Delta t<\frac{1}{2}$, then the maximum point is bigger than $2$, so since $g(2)=2$, for a sufficiently small $\varepsilon$, $g(x)>2$ if $x\in(2,2+\varepsilon)$. Our goal is to choose $\Delta t$ in a way that $2+\varepsilon>\tilde{y}$.
    If we solve $g(x)=2$, the solution (apart from the previously mentioned $x_{0}=2$) is $x_{1}=\frac{1}{\Delta t}$. Consequently, if $\Delta t\leq\frac{1}{\tilde{y}}$, then the scheme preserves the lower bound of the method. ∎
In conclusion, a proper choice for parameter $\mathcal{B}$ is $\mathcal{B}:=\mathcal{C}\min\left\{\dfrac{1}{2},\dfrac{1}{\tilde{y}}\right\}$, and function $\varphi$ can be constructed in accordance with the ideas mentioned in Section 3.3. Therefore, in line with the results of Theorems 3 and 4, we expect the numerical solution produced by the nonstandard multistep method to be bounded and attain property (11) for all $\Delta t\geq 0$ step sizes. In this case, Theorem 5 is not applicable since the equation is one-dimensional (and the preservation of the constant values is trivial).
In the following, we run several numerical tests observing the performance of the aforementioned numerical methods. For this, we use the strongly stable forms of multistep methods [52] - the coefficients of the methods applied are listed in the Appendix. Moreover, since the solution is known, we used the exact values for those initial values of the numerical method that are needed (for example, for an $s$-step method, one needs the values of $u^{0},\;u_{1},\;\dots,\;u^{s-1}$). Alternatively, we could have used some further, high-order (possibly nonstandard) numerical method to approximate these values - such an approach is taken in the second example in Section 5.2.
[…]
$\Delta t$ $\varphi_{1}$ errors $\varphi_{1}$ orders $\varphi_{2}$ errors $\varphi_{2}$ orders $\varphi_{3}$ errors $\varphi_{3}$ orders
$0.1$ $1.4009\cdot 10^{-1}$ $1.1611\cdot 10^{-1}$ $1.9461\cdot 10^{-1}$
$0.1\cdot 2^{-1}$ $1.0611\cdot 10^{-1}$ $0.4008$ $8.2200\cdot 10^{-2}$ $0.4983$ $1.7290\cdot 10^{-1}$ $0.1706$
$0.1\cdot 2^{-2}$ $5.8780\cdot 10^{-2}$ $0.8521$ $4.44178\cdot 10^{-2}$ $0.8958$ $1.0622\cdot 10^{-1}$ $0.7028$
$0.1\cdot 2^{-3}$ $3.0750\cdot 10^{-2}$ $0.9347$ $2.2833\cdot 10^{-2}$ $0.9522$ $5.8599\cdot 10^{-2}$ $0.8582$
$0.1\cdot 2^{-4}$ $1.5669\cdot 10^{-2}$ $0.9727$ $1.1576\cdot 10^{-2}$ $0.9799$ $3.0621\cdot 10^{-2}$ $0.9364$
$0.1\cdot 2^{-5}$ $7.9013\cdot 10^{-3}$ $0.9877$ $5.8249\cdot 10^{-3}$ $0.9909$ $1.5626\cdot 10^{-2}$ $0.9706$
$0.1\cdot 2^{-6}$ $3.9666\cdot 10^{-3}$ $0.9942$ $2.9212\cdot 10^{-3}$ $0.9957$ $7.8894\cdot 10^{-3}$ $0.9860$
$0.1\cdot 2^{-7}$ $1.9871\cdot 10^{-3}$ $0.9972$ $1.4627\cdot 10^{-3}$ $0.9979$ $3.9634\cdot 10^{-3}$ $0.9932$
$0.1\cdot 2^{-8}$ $9.9452\cdot 10^{-4}$ $0.9986$ $7.3190\cdot 10^{-4}$ $0.9990$ $1.9863\cdot 10^{-3}$ $0.9966$
$0.1\cdot 2^{-9}$ $4.9750\cdot 10^{-4}$ $0.9993$ $3.6608\cdot 10^{-4}$ $0.9995$ $9.9431\cdot 10^{-4}$ $0.9983$
$\Delta t$ $\varphi_{4}$ errors $\varphi_{4}$ orders $\varphi_{5}$ errors $\varphi_{5}$ orders $\varphi_{6}$ errors $\varphi_{6}$ orders
$0.1$ $1.4359\cdot 10^{-1}$ $9.7188\cdot 10^{-2}$ $1.1764\cdot 10^{-1}$
$0.1\cdot 2^{-1}$ $8.2705\cdot 10^{-2}$ $0.7959$ $4.1449\cdot 10^{-2}$ $1.2294$ $5.7578\cdot 10^{-2}$ $1.0308$
$0.1\cdot 2^{-2}$ $2.7204\cdot 10^{-2}$ $1.6041$ $1.1739\cdot 10^{-2}$ $1.8200$ $1.7248\cdot 10^{-2}$ $1.7390$
$0.1\cdot 2^{-3}$ $7.5017\cdot 10^{-3}$ $1.8586$ $3.0902\cdot 10^{-3}$ $1.9255$ $4.6113\cdot 10^{-3}$ $1.9032$
$0.1\cdot 2^{-4}$ $1.9405\cdot 10^{-3}$ $1.9508$ $7.8967\cdot 10^{-4}$ $1.9684$ $1.1830\cdot 10^{-3}$ $1.9628$
$0.1\cdot 2^{-5}$ $4.9156\cdot 10^{-4}$ $1.9810$ $1.994\cdot 10^{-4}$ $1.9854$ $2.9904\cdot 10^{-4}$ $1.9840$
$0.1\cdot 2^{-6}$ $1.2358\cdot 10^{-4}$ $1.9919$ $5.0099\cdot 10^{-5}$ $1.9930$ $7.5143\cdot 10^{-5}$ $1.9926$
$0.1\cdot 2^{-7}$ $3.0976\cdot 10^{-5}$ $1.9963$ $1.2555\cdot 10^{-5}$ $1.9966$ $1.8832\cdot 10^{-5}$ $1.9965$
$0.1\cdot 2^{-8}$ $7.7534\cdot 10^{-6}$ $1.9982$ $3.1424\cdot 10^{-6}$ $1.9983$ $4.7135\cdot 10^{-6}$ $1.9983$
$0.1\cdot 2^{-9}$ $1.9395\cdot 10^{-6}$ $1.9991$ $7.8606\cdot 10^{-7}$ $1.9992$ $1.1791\cdot 10^{-6}$ $1.9991$
$\Delta t$ $\varphi_{7}$ errors $\varphi_{7}$ orders $\varphi_{8}$ errors $\varphi_{8}$ orders
$0.1$ $8.9836\cdot 10^{-2}$ $7.6103\cdot 10^{-2}$
$0.1\cdot 2^{-1}$ $2.4513\cdot 10^{-2}$ $1.8738$ $1.1542\cdot 10^{-2}$ $2.7211$
$0.1\cdot 2^{-2}$ $3.5736\cdot 10^{-3}$ $2.7781$ $8.1974\cdot 10^{-4}$ $3.8155$
$0.1\cdot 2^{-3}$ $4.6978\cdot 10^{-4}$ $2.9273$ $5.3510\cdot 10^{-5}$ $3.9373$
$0.1\cdot 2^{-4}$ $5.9937\cdot 10^{-5}$ $2.9705$ $3.4099\cdot 10^{-6}$ $3.9720$
$0.1\cdot 2^{-5}$ $7.5646\cdot 10^{-6}$ $2.9861$ $2.1515\cdot 10^{-7}$ $3.9863$
$0.1\cdot 2^{-6}$ $9.5005\cdot 10^{-7}$ $2.9932$ $1.3511\cdot 10^{-8}$ $3.9932$
$0.1\cdot 2^{-7}$ $1.1904\cdot 10^{-7}$ $2.9966$ $8.4697\cdot 10^{-10}$ $3.9956$
$0.1\cdot 2^{-8}$ $1.4898\cdot 10^{-8}$ $2.9982$ $5.4143\cdot 10^{-11}$ $3.9675$
$0.1\cdot 2^{-9}$ $1.8655\cdot 10^{-9}$ $2.9974$ $5.673\cdot 10^{-12}$ $3.2545$
Table 1: The errors and the corresponding orders of the different methods using different $\varphi$ functions.
Figure 1: The errors of the different methods with different $\varphi$ functions for different values of $\Delta t$.
[…]
One surprising property is the remarkable initial speed of the convergence of the third-order method, but its order decreases to the expected value of $3$ as the timestep becomes smaller. This might be the result of the effect of the function $\varphi_{8}(\Delta t)$, since this function enables higher order (see the case of SSPMS(6,4)), but this order is decreased when $\Delta t$ is smaller (and where the difference of function $\varphi_{8}(\Delta t)$ from the usual $\varphi(\Delta t)=\Delta t$ choice is less significant). This might indicate that for carefully chosen $\varphi(\Delta t)$ functions, nonstandard multistep methods might have larger orders than their standard counterparts. Another possible reason for this is the fact that we used the exact values as the initial $s$-many values of the method - if we use only a third-order approximation for these, one can only get third-order convergence (as we can see in Section 5.2).
Figure 2: The errors of the different multistep methods for different values of $\Delta t$.
NSSPMS(4,2) NSSPMS(4,3) NSSPMS(6,4)
$\Delta t$ errors orders errors orders errors orders
$0.05$ $1.6660\cdot 10^{-4}$ $8.2145\cdot 10^{-4}$ $1.1541\cdot 10^{-2}$
$0.05\cdot 2^{-1}$ $6.0870\cdot 10^{-5}$ $1.4527$ $5.7502\cdot 10^{-5}$ $3.8365$ $8.1974\cdot 10^{-4}$ $3.8155$
$0.05\cdot 2^{-2}$ $1.7144\cdot 10^{-5}$ $1.8280$ $4.1033\cdot 10^{-6}$ $3.8087$ $5.3510\cdot 10^{-6}$ $3.9373$
$0.05\cdot 2^{-3}$ $4.4918\cdot 10^{-6}$ $1.9324$ $3.1262\cdot 10^{-7}$ $3.7143$ $3.4099\cdot 10^{-6}$ $3.9720$
$0.05\cdot 2^{-4}$ $1.1463\cdot 10^{-6}$ $1.9703$ $2.6326\cdot 10^{-8}$ $3.5698$ $2.1515\cdot 10^{-7}$ $3.9863$
$0.05\cdot 2^{-5}$ $2.8934\cdot 10^{-7}$ $1.9862$ $2.4865\cdot 10^{-9}$ $3.4043$ $1.3511\cdot 10^{-8}$ $3.9932$
$0.05\cdot 2^{-6}$ $7.2670\cdot 10^{-8}$ $1.9933$ $2.6035\cdot 10^{-10}$ $3.2556$ $8.4697\cdot 10^{-10}$ $3.9956$
$0.05\cdot 2^{-7}$ $1.8208\cdot 10^{-8}$ $1.9967$ $2.9433\cdot 10^{-11}$ $3.1450$ $5.4143\cdot 10^{-11}$ $3.9675$
$0.05\cdot 2^{-8}$ $4.5571\cdot 10^{-9}$ $1.9984$ $3.584\cdot 10^{-12}$ $3.0376$ $5.673\cdot 10^{-12}$ $3.2545$
Table 2: The errors and the corresponding orders of the different multistep methods.
Finally, we observe the way the nonstandard methods preserve the qualitative properties of the original differential equation. By Theorems 3 and 4, we know that the nonstandard multistep method preserves the boundedness property of the original solution (since the Forward Euler method has this property for $\Delta t\leq\mathcal{B}$, see Proposition 1), and condition (11) also holds. In Figure 3, we compare the results of the nonstandard methods with their standard counterparts applied with the same step sizes, initial value $\tilde{y}=3$, and function choices $\varphi_{5}$, $\varphi_{7}$, and $\varphi_{8}$, respectively. As it can be seen, for $\Delta t_{1}=0.5$ the standard methods are highly unstable: they are not bounded from below, and (11) is also violated. However, the nonstandard ones behave as expected, even for a larger timestep: their solutions are bounded from below, and the property (11) also holds with $s=4$ (in the first two cases) and with $s=6$ (in the last case). For $\Delta t_{2}=0.02$, both methods behave as expected.
Figure 3: Methods SSPMS(4,2) and NSSPMS(4,2) (upper left), SSPMS(4,3) and NSSPMS(4,3) (upper right) and SSPMS(6,4) and NSSPMS(6,4) (lower) with timesteps $\Delta t_{1}=0.5$ and $\Delta t_{2}=0.02$. For $\Delta t_{2}$, the schemes are very close to each other, therefore, only one of them is visible.
5.2 The SEIR system
The following form of the Kermack-McKendrick model [55] describes the propagation of an illness among a given constant population:
$S^{\prime}(t)$ $=-5S(t)I(t),$ (17)
$E^{\prime}(t)$ $=5S(t)I(t)-E(t),$
$I^{\prime}(t)$ $=E(t)-I(t),$
$R^{\prime}(t)$ $=I(t),$
where $S(t),E(t),I(t)$ and $R(t)$ describe the fraction of susceptible, exposed (infected but not yet infectious), ill (infected and infectious), and recovered people, respectively. It is easy to see that if $S(t_{0}),E(t_{0}),I(t_{0}),R(t_{0})\geq 0$, then $S(t),E(t),I(t),R(t)\geq 0$ holds for every $t\in[t_{0},T]$. Moreover, $S(t)+E(t)+I(t)+R(t)=\mathcal{M}$ is also true for some constant $\mathcal{M}\in\mathbb{R}^{+}$ for every $t\in[t_{0},T]$ (an interested reader might consult [56]).
Now, let us state a result about the behavior of the forward Euler method, which will be used in the choice of the function $\varphi(x)$.
Proposition 2.
Let us apply the forward Euler method to equation (17) with stepsize $\Delta t_{FE}\leq\min\left\{\dfrac{1}{\mathcal{M}},1\right\}$. Then, the forward Euler method applied to equation (17) preserves the non-negative property of the numerical solution, and $S^{n}+E^{n}+I^{n}+R^{n}=\mathcal{M}$ also holds.
The proof of a more general case can be found in [57]. Consequently, the choice $\mathcal{B}=\mathcal{C}\min\left\{\dfrac{1}{\mathcal{M}},1\right\}$ is appropriate.
A topic that has been omitted until this point is the choice of the initial values of the numerical method. In [49], it was shown that nonstandard versions of Runge-Kutta methods (i.e. ones, where the usual $\Delta t$ term is interchanged with a $\varphi(\Delta t)$ term), have (at least) the same order $p$ as their ’standard’ counterparts and also preserve the linear stability of the system if the following conditions are met:
  • Let us assume that for the ’standard’ numerical method, the linear stability and positivity are preserved if $\Delta t<\varphi^{*}$. Then the first condition is $\varphi(\Delta t)<\varphi^{*}$.
  • $\varphi(\Delta t)=\Delta t+O((\Delta t)^{p+1})$.
Note that both of these conditions are met if we choose a function $\varphi$ by the ideas discussed in Section 3.3 with $\mathcal{B}=\varphi^{*}$. Moreover, in [57] it was shown that ’standard’ Runge-Kutta methods written in the strongly stability preserving form also preserve positivity and the sum of the components if $\Delta t\leq\mathcal{C}\min\left\{\dfrac{1}{\mathcal{M}},1\right\}$ where $\mathcal{C}$ is the SSP coefficient of the Runge-Kutta method. In conclusion, for the initial $11,2,\dots,s-$ 1 values we can use a nonstandard Runge-Kutta method with a suitable order. The forms of the SSP-Runge-Kutta methods used during the experiments are listed in the Appendix.
[…]
Figure 4: The errors for different choices of function $\varphi$.
$\Delta t$ $\varphi_{1}$ errors $\varphi_{1}$ orders $\varphi_{2}$ errors $\varphi_{2}$ orders $\varphi_{3}$ errors $\varphi_{3}$ orders
$0.1$ $1.1531\cdot 10^{-1}$ $8.5437\cdot 10^{-2}$ $2.0912\cdot 10^{-1}$
$0.1\cdot 2^{-1}$ $5.5339\cdot 10^{-2}$ $1.0592$ $4.0419\cdot 10^{-2}$ $1.0798$ $1.0887\cdot 10^{-1}$ $0.9416$
$0.1\cdot 2^{-2}$ $2.6622\cdot 10^{-2}$ $1.0556$ $1.9456\cdot 10^{-2}$ $1.0548$ $5.3511\cdot 10^{-2}$ $1.0247$
$0.1\cdot 2^{-3}$ $1.2993\cdot 10^{-2}$ $1.0349$ $9.5200\cdot 10^{-3}$ $1.0312$ $2.6159\cdot 10^{-2}$ $1.0325$
$0.1\cdot 2^{-4}$ $6.4107\cdot 10^{-3}$ $1.0192$ $4.7058\cdot 10^{-3}$ $1.0165$ $1.2878\cdot 10^{-2}$ $1.0224$
$0.1\cdot 2^{-5}$ $3.1830\cdot 10^{-3}$ $1.0101$ $2.3391\cdot 10^{-3}$ $1.0085$ $6.3822\cdot 10^{-3}$ $1.0128$
$0.1\cdot 2^{-6}$ $1.5859\cdot 10^{-3}$ $1.0051$ $1.1661\cdot 10^{-3}$ $1.0043$ $3.1760\cdot 10^{-3}$ $1.0069$
$0.1\cdot 2^{-7}$ $7.9150\cdot 10^{-4}$ $1.0026$ $5.8217\cdot 10^{-4}$ $1.0022$ $1.5841\cdot 10^{-3}$ $1.0035$
$0.1\cdot 2^{-8}$ $3.9539\cdot 10^{-4}$ $1.0013$ $2.9087\cdot 10^{-4}$ $1.0011$ $7.9106\cdot 10^{-3}$ $1.0018$
$0.1\cdot 2^{-9}$ $1.9761\cdot 10^{-4}$ $1.0007$ $1.4538\cdot 10^{-4}$ $1.0005$ $3.9528\cdot 10^{-4}$ $1.0009$
$\Delta t$ $\varphi_{4}$ errors $\varphi_{4}$ orders $\varphi_{5}$ errors $\varphi_{5}$ orders $\varphi_{6}$ errors $\varphi_{6}$ orders
$0.1$ $8.6047\cdot 10^{-2}$ $3.9627\cdot 10^{-2}$ $5.6965\cdot 10^{-2}$
$0.1\cdot 2^{-1}$ $2.4040\cdot 10^{-2}$ $1.8397$ $1.0037\cdot 10^{-2}$ $1.9811$ $1.4924\cdot 10^{-2}$ $1.9324$
$0.1\cdot 2^{-2}$ $6.1921\cdot 10^{-3}$ $1.9570$ $2.5270\cdot 10^{-3}$ $1.9898$ $3.7831\cdot 10^{-3}$ $1.9800$
$0.1\cdot 2^{-3}$ $1.5642\cdot 10^{-3}$ $1.9850$ $6.3499\cdot 10^{-4}$ $1.9926$ $9.5204\cdot 10^{-4}$ $1.9905$
$0.1\cdot 2^{-4}$ $3.9275\cdot 10^{-4}$ $1.9937$ $1.5924\cdot 10^{-4}$ $1.9955$ $2.3884\cdot 10^{-4}$ $1.9950$
$0.1\cdot 2^{-5}$ $9.8386\cdot 10^{-5}$ $1.9971$ $3.9878\cdot 10^{-5}$ $1.9975$ $5.9816\cdot 10^{-5}$ $1.9974$
$0.1\cdot 2^{-6}$ $2.4620\cdot 10^{-5}$ $1.9986$ $9.9785\cdot 10^{-6}$ $1.9987$ $1.4968\cdot 10^{-5}$ $1.9987$
$0.1\cdot 2^{-7}$ $6.1580\cdot 10^{-6}$ $1.9993$ $2.4958\cdot 10^{-6}$ $1.9993$ $3.7436\cdot 10^{-6}$ $1.9993$
$0.1\cdot 2^{-8}$ $1.5399\cdot 10^{-6}$ $1.9997$ $6.2409\cdot 10^{-7}$ $1.9997$ $9.3613\cdot 10^{-7}$ $1.9997$
$0.1\cdot 2^{-9}$ $3.8502\cdot 10^{-7}$ $1.9998$ $1.5605\cdot 10^{-7}$ $1.9997$ $2.3407\cdot 10^{-7}$ $1.9998$
$\Delta t$ $\varphi_{7}$ errors $\varphi_{7}$ orders $\varphi_{8}$ errors $\varphi_{8}$ orders
$0.1$ $2.2574\cdot 10^{-2}$ $1.0324\cdot 10^{-2}$
$0.1\cdot 2^{-1}$ $3.0024\cdot 10^{-3}$ $2.9105$ $6.8634\cdot 10^{-4}$ $3.9110$
$0.1\cdot 2^{-2}$ $3.8218\cdot 10^{-4}$ $2.9738$ $4.3684\cdot 10^{-5}$ $3.9737$
$0.1\cdot 2^{-3}$ $4.8138\cdot 10^{-5}$ $2.9890$ $2.7538\cdot 10^{-6}$ $3.9876$
$0.1\cdot 2^{-4}$ $6.0390\cdot 10^{-6}$ $2.9948$ $1.7287\cdot 10^{-7}$ $3.9937$
$0.1\cdot 2^{-5}$ $7.5622\cdot 10^{-7}$ $2.9974$ $1.0829\cdot 10^{-8}$ $3.9967$
$0.1\cdot 2^{-6}$ $9.4613\cdot 10^{-8}$ $2.9987$ $6.7923\cdot 10^{-10}$ $3.9949$
$0.1\cdot 2^{-7}$ $1.1835\cdot 10^{-8}$ $2.9990$ $4.5636\cdot 10^{-11}$ $3.8957$
$0.1\cdot 2^{-8}$ $1.4857\cdot 10^{-9}$ $2.9939$ $9.0434\cdot 10^{-12}$ $2.3352$
$0.1\cdot 2^{-9}$ $1.9758\cdot 10^{-10}$ $2.9106$ $1.2814\cdot 10^{-11}$
Table 3: The errors and the corresponding orders of the different methods using different $\varphi$ functions.
[…]
Figure 5: The errors of the different multistep methods for different values of $\Delta t$.
NSSPMS(4,2) NSSPMS(4,3) NSSPMS(6,4)
$\Delta t$ errors orders errors orders errors orders
$0.05$ $3.7854\cdot 10^{-4}$ $4.9316\cdot 10^{-4}$ $6.1934\cdot 10^{-4}$
$0.05\cdot 2^{-1}$ $1.0095\cdot 10^{-4}$ $1.9068$ $9.1717\cdot 10^{-6}$ $2.4267$ $4.2076\cdot 10^{-5}$ $3.8797$
$0.05\cdot 2^{-2}$ $2.6162\cdot 10^{-5}$ $1.9481$ $1.4250\cdot 10^{-6}$ $2.6862$ $2.7473\cdot 10^{-6}$ $3.9369$
$0.05\cdot 2^{-3}$ $6.6660\cdot 10^{-6}$ $1.9726$ $1.9672\cdot 10^{-7}$ $2.8568$ $1.7543\cdot 10^{-7}$ $3.9690$
$0.05\cdot 2^{-4}$ $1.6829\cdot 10^{-6}$ $1.9859$ $2.5792\cdot 10^{-8}$ $2.9312$ $1.1084\cdot 10^{-8}$ $3.9845$
$0.05\cdot 2^{-5}$ $4.2280\cdot 10^{-7}$ $1.9928$ $3.3003\cdot 10^{-9}$ $2.9662$ $6.9645\cdot 10^{-10}$ $3.9923$
$0.05\cdot 2^{-6}$ $1.0597\cdot 10^{-7}$ $1.9964$ $4.1735\cdot 10^{-10}$ $2.9833$ $4.3562\cdot 10^{-11}$ $3.9989$
$0.05\cdot 2^{-7}$ $2.6525\cdot 10^{-8}$ $1.9982$ $5.2470\cdot 10^{-11}$ $2.9917$ $2.5479\cdot 10^{-12}$ $4.0957$
$0.05\cdot 2^{-8}$ $6.6354\cdot 10^{-9}$ $1.9991$ $6.5846\cdot 10^{-12}$ $2.9943$ $1.1435\cdot 10^{-12}$ $1.1558$
Table 4: The errors and the corresponding orders of the different multistep methods.
For the last test concerning the SEIR model, we compare the behavior of the nonstandard methods and the standard counterparts for two different timesteps (the unmentioned parameters remain the same) for $t\in[0,10]$. Here we used Runge-Kutta methods again with an appropriate order to approximate the initial values of the scheme: standard ones for the standard multistep methods and nonstandard ones (with appropriate choices of $\varphi$) for the nonstandard versions. In all three cases, we only plot the numerical solutions of $I^{n}$, but the others also show signs of similar behavior. In Figure 6, we can see that the standard methods behave in an unstable way: they become negative, while the nonstandard ones remain in the upper half plane for the whole process, even for larger timesteps. Also, for a small timestep, the solutions are very close to each other and both of them behave as expected.
Figure 6: The comparison of the standard and nonstandard methods. Upper left: SSPMS(4,2) and NSSPMS(4,2) with $\Delta t_{1}=10/12$ and $\Delta t_{2}=0.02$. Upper right: SSPMS(4,3) and NSSPMS(4,3) with $\Delta t_{1}=10/15$ and $\Delta t_{2}=0.02$. Lower: SSPMS(6,4) and NSSPMS(6,4) with $\Delta t_{1}=10/15$ and $\Delta t_{2}=0.02$. The solutions for $\Delta t_{2}$ time-step are very close to each other in all three cases.
As a final remark, we mention that because of the form of the numerical schemes, the sum $S^{n}+E^{n}+I^{n}+R^{n}$ is preserved for every $\Delta t>0$ for the standard and nonstandard methods alike, so such numerical tests are omitted from the paper.
6 Conclusions and discussion
In this work, we considered nonstandard multistep methods. By using a nonstandard version of the classical Taylor series (Lemma 1), we were able to show that (under some conditions for the denominator function $\varphi$) the nonstandard versions of these methods have the same order as their standard counterparts. Note that Lemma 1 might also be useful in the observation of other, high-order nonstandard methods that are constructed in a similar fashion, i.e., involving a generalized denominator function.
In Section 3.2, it was shown that (under some further conditions for function $\varphi$) the nonstandard methods preserve the boundedness (and as a special case, positivity) properties of the original, continuous problem. To achieve this goal, a bound for the timestep should be calculated under which the Euler method preserves these properties. We note that the elementary stable property was not considered, but theorems similar to those in [49] might be proven in the future for nonstandard multistep methods too. These results would involve the observation of the elementary stability property of standard multistep methods, which might turn out to be harder (in general, these are not even true [50], as mentioned in Section 2) - this could also be a topic for further research. On the other hand, the observation of forward Euler methods might turn out to be easier, and for some systems, the preservation of positivity ensures elementary stability (see, for example, the case of an integro-differential equation [58] and its delayed version [59]).
In Section 3.3 we also presented several possible choices for the function $\varphi$, for which the conditions of the theorems of the previous sections are met. We also mentioned there that in [49], the authors gave a different construction of similar functions - in a further work, one could also compare the performance of the functions of this article and those defined in [49].
In Section 5, we performed some tests for two test problems and observed that the nonstandard linear multistep methods behave as expected. In further articles, the same methods could also be applied to more robust ordinary systems and also to partial differential equations. One could also compare the performance of these multistep methods with other, higher-order methods (for example, the ones mentioned in Section 1).
Another interesting fact was the high order of convergence in the case of the method NSSPMS(4,3) in Section 5.1, which might have been caused by the use of the exact solutions for the initial values of the scheme. A further investigation in this regard (e.g., the increase of the rate of convergence in the case of higher order approximations of the initial values) might also prove to be fruitful. Finally, it would also be beneficial to observe the conditions under which some nonstandard multistep methods are A-stable and check whether the barriers of Dahlquist (see e.g. [51]) also hold in these nonstandard cases.
Acknowledgement
The author would like to thank Prof. István Faragó for his insightful comments during the preparation of the article. This research has been supported by the National Research, Development and Innovation Office – NKFIH, grant no. K137699. The research reported in this paper is also part of project no. BME-NVA-02, implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021 funding scheme.
Declarations
Conflict of interest The author declares that they have no conflict of interest.
Appendix A The coefficients of the different methods
Here we list the coefficients of the different strongly-stable multistep and Runge-Kutta methods used in the paper, along with their SSP coefficients.
  • Explicit SSPMS(4,2) [60], with $\mathcal{C}=\frac{2}{3}$:
    $\tilde{\alpha}_{1}=\dfrac{8}{9},\quad\tilde{\beta}_{1}=\dfrac{4}{3},\quad\tilde{\alpha}_{4}=\dfrac{1}{9}.$
  • Explicit SSPMS(4,3) [60], with $\mathcal{C}=\frac{1}{3}$:
    $\tilde{\alpha}_{1}=\dfrac{16}{27},\quad\tilde{\beta}_{1}=\dfrac{16}{81},\quad\tilde{\alpha}_{4}=\dfrac{11}{27},\quad\tilde{\beta}_{4}=\dfrac{4}{9}.$
  • Explicit SSPMS(6,4) [61, 62], with $\mathcal{C}=0.1648$:
[…]
  • Explicit SSPRK(2,2) with $\mathcal{C}=1$ [63]:
    $u^{(1)}$ $=u^{n}+\Delta tf(u^{n}),$
    $u^{n+1}$ $=\dfrac{1}{2}u^{n}+\dfrac{1}{2}\left(u^{(1)}+\Delta tf(u^{(1)})\right).$
  • Explicit SSPRK(3,3) with $\mathcal{C}=1$ [52]:
    $u^{(1)}$ $=u^{n}+\Delta tf(u^{n}),$
    $u^{(2)}$ $=\dfrac{3}{4}u^{n}+\dfrac{1}{4}u^{(1)}+\dfrac{1}{4}\Delta tf(u^{(1)}),$
    $u^{n+1}$ $=\dfrac{1}{3}u^{n}+\dfrac{2}{3}\left(u^{(2)}+\Delta tf(u^{(2)})\right).$
  • Explicit SSPRK(4,10) with $\mathcal{C}=6$ [62]:
    $u^{(1)}$ $=u^{n}+\dfrac{1}{6}\Delta tf(u^{n}),$
    $u^{(j+1)}$ $=u^{(j)}+\dfrac{1}{6}\Delta tf(u^{(j)}),\qquad j=1,2,3,$
    $u^{(5)}$ $=\dfrac{3}{5}u^{n}+\dfrac{2}{5}u^{(4)}+\dfrac{1}{15}\Delta tf(u^{(4)}),$
    $u^{(j+1)}$ $=u^{(j)}+\dfrac{1}{6}\Delta tf(u^{(j)}),\qquad j=5,\dots,8,$
    $u^{n+1}$ $=\dfrac{1}{25}u^{n}+\dfrac{9}{25}u^{(4)}+\dfrac{3}{5}u^{(9)}+\dfrac{3}{50}\Delta tf(u^{(4)})+\dfrac{1}{10}\Delta tf(u^{(9)}).$